# Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting

### v2

@@ Abstract
Time series forecasting task predicts future trends based on historical information. Transformer-based U-Net architectures, despite their success in medical image segmentation, have limitations in both expressiveness and computation efficiency in time series forecasting as evidenced in YFormer. To tackle these challenges, we introduce Kernel-U-Net, a symmetric and hierarchical U-shape neural network architecture. The kernel-U-Net encoder compresses gradually input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series. Specifically, Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear. Experiments on seven real-world datasets, considering both multivariate and univariate settings, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model PatchTST in the majority of cases and outperforms Yformer. The source code for Kernel-U-Net will be made publicly available for further research and application.
1 Introduction
@@
Time series forecasting predicts future trends based on recent historical information. It allows experts to track the incoming situation and react timely in critical cases. Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city Lai et al. (2018), monitoring influenza-like illness weekly patient cases Wu et al. (2021), monitoring electricity transformer temperature in the electric power long-term deployment Zhou et al. (2021) or forecasting temperature, pressure and humidity in weather station Liu et al. (2022b) etc.
@@
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methodsKhandelwal et al. (2015) and machine learning techniquesPersson et al. (2017) to deep learning-based solutions, such as recurrent neural networks(RNN) Tokgöz and Ünal (2018), Long Short-term Memory (LSTM) Kong et al. (2017), Temporal Convolutional Network (TCN) Hewage et al. (2020) and Transformer-based model Li et al. (2019).
@@
Among the Transformer models applying to time series data, Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022) are the best variants that incrementally improved the quality of prediction. As a recent paper Zeng et al. (2023) challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in Nie et al. (2023) argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches. By adding a linear patch layer, their model PatchTST successfully relieved the overfitting problem of transformer modules and reached state-of-the-art results.
@@
The Convolutional U-net, as a classic and highly expressive model in medical image segmentationRonneberger et al. (2015), features a symmetric encoder and decoder structure that is elegant in its design. This model’s structure is particularly suited to the time series forecasting task, as both inputs and outputs in this context are typically derived from the same distribution. The first U-shape model adapted for time series forecasting was the YformerMadhusudhanan et al. (2023), which incorporated Transformers in both its encoder and decoder components. As mentioned previously, employing Transformers on point-wise data has the potential to cause overfitting issues. Therefore, our investigation aims to discover if there is a U-shape architecture effective in time series forecasting, which also possesses the capability to integrate various modules flexibly, thus facilitating the specific customization of solutions.
To tackle this problem, we propose a symmetric and hierarchical architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net and Yformer for time series forecasting. K-U-Net provides convenience for composing particular models with custom kernels that follow the design pattern. By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data. Furthermore, the hierarchical structure of K-U-Net exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning such sequences. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
To fully study the performance and efficiency of K-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. We compose 30 variants of K-U-Net by placing different kernels at different layers and then we choose the best candidates for fine-tuning. Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinearZeng et al. (2023) and PatchTST, in the majority of cases.
@@
  • We propose Kernel-U-Net, a symmetric and hierarchical architecture that partitions input sequences into small patches at each layer of the network.
  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the design pattern.
@@
We conclude that Kernel-U-Net stands out as a highly promising option for large-scale time series forecasting tasks. Its symmetric and hierarchical design provides a balance of low computational complexity and high expressiveness. In most scenarios, it either surpasses or is slightly below the state-of-the-art results. Furthermore, its adaptability in fast-paced development and testing environments is ensured by the use of flexible, customizable kernels.
Figure 1: Comparison of the structures of different U-Net. a) classic Convolutional U-Net, b) Kernel U-Net, c) Swin Transformer U-Net
2 Related works
Transformer
@@ 2 Related works
Transformer Vaswani et al. (2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks. It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route.
@@
Vision Transformers (ViTs) Dosovitskiy et al. (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al. (2009). Swin Transformer Liu et al. (2021) proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
@@
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans Li et al. (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. To improve the computation efficiency, the authors propose also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer Zhou et al. (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. The authors propose a ProbSparse self-attention mechanism to replace the canonical self-attention efficiently. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. PyraformerLiu et al. (2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass. PatchTST employed PropSparse and patches, thus reducing computation complexity to $O(\frac{L}{S}log(\frac{L}{S}))$, where $S$ is the size of the patch.
@@
Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing. Autoformer Wu et al. (2021) introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods. Frequency Enhanced Decomposed Transformer (FEDformer)Zhou et al. (2022) captures global properties of time series with seasonal-trend decomposition. The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
U-Net
@@
U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al. (2015). U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is reduced by a convolutional layer and a max-pooling layer gradually into a latent vector. At the decoder stage, the latent vector is developed by a transposed convolutional layer for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
@@
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al. (2015), biomedical 3D-Image segmentation Çiçek et al. (2016), time series segmentation Perslev et al. (2019), image super-resolution Han et al. (2022) and image denoising Zhang et al. (2023). Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al. (2016), 1D-U-Net Azar et al. (2022) and Swin-U-Net Cao et al. (2023) that only use transformer blocks at each layer of U-Net.
@@
In time series processing, U-time Perslev et al. (2019) is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer Madhusudhanan et al. (2023) is the first U-Net based model for time series forecasting task. In particular, YFormer applied transformer blocks on each layer of U-Net and capitalized on multi-resolution feature maps effectively.
Hierarchical and hybrid model
@@
In time series processing, the increasing size of data degrades the performance of deep models and increases crucially the cost of learning them. For example, recurrent models such as RNN, LSTM, GRU have a linear complexity but suffer the gradient vanishing problem when input length increases. Transformer-based block captures better long dependencies but requires $O(L^{2})$ computations in general. To balance the expressiveness of complex models and computational efficiency, researchers investigated hybrid models that merge different modules into the network and hierarchical architectures.
@@
For example, authors in Du et al. (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al. (2017) applied hierarchical LSTM and GRU for document classification. In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
@@
To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Duan et al. (2022), pyramidal structure in Pyraformer Liu et al. (2022a), AutoEncoders Asres et al. (2022), U-Net-like model in Yformer Madhusudhanan et al. (2023) or patches in PatchTST Nie et al. (2023).
@@
Kernel-U-Net is a symmetric and hierarchical U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences. Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
3 Method
3.1 Problem Formulation
@@
Let us note by $x\in R^{M\times N}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the size of the feature. Let L be the length of memory or the look-back window, so $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$ is a slice of length $L$ of all features. It contains historical information about the system at instant $t$. We call $X\in R^{N\times M\times L}$ the trajectory matrix which represents the sliced historical data for each instant $t\in[0,N-L-1]$.
@@
In the context of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$. Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window. Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps. Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
3.2 Kernel U-Net
Kernel U-Net(K-U-Net) is a hierarchical and symmetric U-shape architecture. It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels(Figure 1). More precisely, the K-U-Net encoder reshapes the input sequences into a large batch of small patches and repeatedly applies custom kernels on them until obtaining the latent vectors. Later, the K-U-Net decoder develops the batch of latent vectors into patches gradually at each layer and obtains a large batch of small patches(Figure 2). At last, K-U-Net reshapes the patches to get the final output. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
Figure 2: Hierarchical Partition
Algorithm 1 Kernel Wrapper
Input: kernel $f_{i}$, input patches $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
Output: Instance of Kernel Wrapper
  # Define the init function :
  def __init(inputs)__:
  # Initiation of main kernel function $f$
  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
  # Holds variables for global operation
  _unet_skip_output = None
  _unet_skip_input = None
  # Define the forward function :
  def forward(x):
  reshape $x$ into (-1, $L_{in}$,$M_{in}$)
  if _unet_skip_input is not None then
           $x=x+\_unet\_skip\_input$
  end if
  # Apply $f$ on $x$
  $y=f(x)$
  _unet_skip_output = $y$
  return $y$
Hierarchical Partition of the input sequence.
Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step. Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small patches $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all patches is equal to its original trajectory matrix:
$X_{t}=Concat([[X_{t,i,j}\forall i\in[1,\frac{L}{L_{1}}]]\forall j\in[1,\frac{M}{M_{1}}]])$.
The partitioned sequences form a large batch of patches and will be processed by kernels in the encoder. Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated patches $\hat{X}_{t+L,i,j}$.
Hierarchical processing with kernels in the encoder and decoder
Let us consider $X$ as a batch of trajectory matrix $X\in R^{B\times L\times M}$, where $B$ is the batch size. The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage. By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder.
At the stage of the encoder, the shape of input $X$ is $(B,L,M)$. Let us note $L_{1}$ and $M_{1}$ as the length and feature size of a patch. Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem. We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of patches and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$. After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer. Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector (Algorithm 2).
At the stage of the decoder, the operations are reversed. We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels. At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$ (Algorithm 3)..
The algorithm is optimized in implementation because the for loop is expensive in the forward function of a PyTorch module. In our implementation, we nest the reshape and transpose operations in the kernel wrapper to avoid iteration over kernels. The complete algorithm is described in the next section.
Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns. A kernel wrapper inherits ”nn.Module” class in Pytorch and offers an interface that executes the given kernel inside the encoder and decoder. It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector (Algorithm 1). By default, a kernel receives a batch of patches in the shape of $(-1,L_{i},M_{i})$ and outputs a new batch of patches in the shape of $(-1,L_{i+1},M_{i+1})$. Remark that the kernels reduce the length of patches from $L_{i}$ to $1$ in the encoder and expand the length of patches from $1$ to $L_{i}$ in the decoder. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function. It reshapes the input patches and checks the output shape outside of the kernel.
Algorithm 2 Kernel U-Net Encoder
Input: $L_{1},M_{1}$, $[L_{2},$,...,… ,$L_{n},$ $L_{n+2},$,...,… ,$L_{n+m}]$, $[M_{2},$,...,… ,$M_{n},$ $M_{n+2},$,...,… ,$M_{n+m}]$, $[f_{1},f_{2},$,...,… ,$f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
@@ 3.2 Kernel U-Net
           layers.append(KernelWrapper($f_{i},$ $L_{i},$ $M_{i},$ $1,$
           $M_{i+1}$))

### v3

# Kernel-U-Net: Multivariate Time Series Forecasting using Custom Kernels

@@ Abstract
Time series forecasting task predicts future trends based on historical information. Transformer-based U-Net architectures, despite their success in medical image segmentation, have limitations in both expressiveness and computation efficiency in time series forecasting as evidenced in YFormer. To tackle these challenges, we introduce Kernel-U-Net, a flexible and kernel-customizable U-shape neural network architecture. The kernel-U-Net encoder compresses the input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series. Specifically, Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; and 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear. Experiments on seven real-world datasets, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model in the majority of cases in channel-independent settings. The source code for Kernel-U-Net will be made publicly available for further research and application.
publicationid: pubid: 979-8-3503-6813- 0/24/\$31.00 ©2024 IEEE
1 Laboratoire Images, Signaux et Systèmes Intelligents (LISSI), Université Paris-Est Créteil , Île-de-France, France
2 Département Informatique et Télécommunication, ESIEE Paris-Université Gustave Eiffel, Île-de-France, France
3 HN-Services, Île-de-France, France
4 Artificial Intelligence Laboratory, UMT, Tirana-Albanie
Proceedings of the 18th International Conference On Innovations In Intelligent Systems And Applications, Craiova, Romania, 2024.
I Introduction
@@
Time series forecasting predicts future trends based on recent historical information. It allows experts to track the incoming situation and react timely in critical cases. Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city [1], monitoring influenza-like illness weekly patient cases [2], monitoring electricity transformer temperature in the electric power long-term deployment [3] or forecasting temperature, pressure and humidity in weather station [4] etc.
Figure 1: Illustration of Kernel U-Net Architecture. a) Architecture of Kernel U-Net, it allows executing linear kernel and nonlinear kernels such as MLP, LSTM, and Transformer. b) Illustration of K-U-Net Encoder, the application of $\phi^{(l)}_{enc}$ on patches is independent of the choice of kernel. c) In the K-U-Net decoder, a custom kernel $\phi^{(l)}_{dec}$ expands the vectors into patches in the reverse order.
@@
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods[5] and machine learning techniques[6] to deep learning-based solutions, such as recurrent neural networks(RNN) [7], Long Short-term Memory (LSTM) [8], Temporal Convolutional Network (TCN) [9] and Transformer-based model [10].
@@
Among the Transformer models applying to time series data, Informer [3], Autoformer [2], and FEDformer [11] are the best variants that incrementally improved the quality of prediction. As a recent paper [12] challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in [13] argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches. By adding a linear patch layer, their model PatchTST successfully relieved the overfitting problem of transformer modules and reached state-of-the-art results.
@@
The Convolutional U-net, as a classic and highly expressive model in medical image segmentation[14], features a symmetric encoder and decoder structure that is elegant in its design. This model’s structure is particularly suited to the time series forecasting task, as both inputs and outputs in this context are typically derived from the same distribution. The first U-shape model adapted for time series forecasting was the Yformer[15], which incorporated Transformers in both its encoder and decoder components. As mentioned previously, employing Transformers on point-wise data has the potential to cause overfitting issues. Therefore, our investigation aims to discover if there is a U-shape architecture effective in time series forecasting, which also possesses the capability to integrate various modules flexibly, thus facilitating the specific customization of solutions.
To tackle this problem, we propose a flexible and kernel-customizable architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net, Swin Transformer, and Yformer for time series forecasting. K-U-Net generalizes the concept of convolutional kernel and provides convenience for composing particular models with non-linear kernels. Following the design pattern, K-U-Net can easily integrate custom kernels by replacing linear kernels with Transformer or LSTM kernels. As a result, K-U-Net can gain expressivity by capturing more complex patterns and dependencies in the data.
Furthermore, the hierarchical structure of K-U-Net exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning such sequences. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
To fully study the performance and efficiency of K-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. We compose 30 variants of K-U-Net by placing different kernels at different layers and then we choose the best candidates for fine-tuning. Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinear[12] and PatchTST, in the majority of cases.
@@
  • We propose Kernel-U-Net, a U-shape architecture that progressively compresses the input sequence into a latent vector and expands it to generate the output sequence.
  • Kernel-U-Net generalizes the concept of the convolutional kernel and provides convenience for composing particular models with custom kernels.
@@
We conclude that Kernel-U-Net stands out as a highly promising option for large-scale time series forecasting tasks. Its hierarchical design provides a balance of low computational complexity and high expressiveness. In most scenarios, it either surpasses or is slightly below the state-of-the-art results. Furthermore, its adaptability in fast-paced development and testing environments is ensured by the use of flexible, customizable kernels.
II Related works
II-1 Transformer
@@
Transformer [16] was initially introduced in the field of Natural Language Processing (NLP) on language translation tasks. It contains a layer of positional encoding, blocks composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route.
@@
Vision Transformers (ViTs) [17] applied a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNet[18]. Swin Transformer [19] proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
@@
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans [10] proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. To improve the computation efficiency, the authors propose also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer [3] has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. The authors propose a ProbSparse self-attention mechanism to replace the canonical self-attention efficiently. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. Pyraformer[20] simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass. PatchTST employs ProbSparse and a linear projection patch layer to reduce the computational complexity of the transformer to $O(\frac{L}{S}\log(\frac{L}{S}))$, where $S$ is patch size. However, its flattened head still incurs a computational cost of $O(\frac{L^{2}}{S})$.
@@
Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing. Autoformer [2] introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods. Frequency Enhanced Decomposed Transformer (FEDformer)[11] captures global properties of time series with seasonal-trend decomposition. The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
II-2 U-Net
@@
U-Net is a neural network architecture designed primarily for medical image segmentation [14]. U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is gradually reduced by a convolutional layer and a max-pooling layer into a latent vector. At the decoder stage, the latent vector is developed by a transposed convolutional layer for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
@@
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation [14], biomedical 3D-Image segmentation [21], time series segmentation [22], image super-resolution [23] and image denoising [24]. Its techniques evolve from basic 2D-U-Net to 3D U-Net [21], 1D-U-Net [25] and Swin-U-Net [26] that only use transformer blocks at each layer of U-Net.
@@
In time series processing, U-time [22] is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer [15] is the first U-Net based model for time series forecasting task. In particular, YFormer applied transformer blocks on each layer of U-Net and capitalized on multi-resolution feature maps effectively.
II-3 Hierarchical and hybrid model
@@
In time series processing, the increasing size of data degrades the performance of deep models and, crucially, increases the cost of learning them. For example, recurrent models such as RNN, LSTM, GRU have a linear complexity but suffer the gradient vanishing problem when input length increases. Transformer-based block captures better long dependencies but requires $O(L^{2})$ computations in general. To balance the expressiveness of complex models and computational efficiency, researchers investigated hybrid models that merge different modules into the network and hierarchical architectures.
@@
For example, authors in [27] investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in [28] stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in [29] applied hierarchical LSTM and GRU for document classification. In time series processing, the authors in [30] combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
@@
To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Swin-Transformer [19], pyramidal structure in Pyraformer [20], U-shape structure in Yformer [15] or patch layer in PatchTST [13].
@@
Kernel-U-Net is a U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences. Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
III Method
III-A Problem Formulation
@@
Let us note by $x\in\mathbb{R}^{N\times M}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the feature size. Let $L$ be the length of memory or the look-back window, we denote the historical time series $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$. We also denote the future time series $(x_{t+L+1},...,x_{t+L+T})$, where $T$ is the length of future horizon and $t\in[0,N-L-T]$ is the time stamp.
@@
The time series forecasting task takes a multivariate time series as input and predicts a future series. Let $x_{t}$ be the features at the time step $t$, $L$ the length of the look-back window, and $T$ the future horizon. Given a historical data series $(x_{t+1},...,x_{t+L})$, time series forecasting task predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future. Then we can define the basic time series forecasting problem:
$(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ (1)
where $f$ is the function that predicts the future series based on a historical series.
Algorithm A Kernel U-Net Encoder
Input: $L_{1}$, $M_{1}$, $\{L_{2}$,...,$L_{n}$,$M_{2}$,...,$M_{m}\}$, $\{D_{2}$,...,$D_{n}$, $D_{n+2}$,..., $D_{n+m-1}\}$, $\{\phi^{(1)}$,...,$\phi^{(n)}$, $\phi^{(n+2)}$,...,$\phi^{(n+m)}\}$, $J_{h}$, $D_{h}$
@@
           layers.append(KW($\phi^{(l)},$ $J_{l},$ $D_{l},$ $1,$ $D_{l\_next}$))